### v1

# GMODiff: One-Step Gain Map Refinement with Diffusion Priors for HDR Reconstruction

Abstract
[…]
1 Introduction
Figure 1: Compared to DNN-based methods, LDM-based method generates perceptually compelling HDR-like images, demonstrating their potential for HDR reconstruction. However, they suffer from high computational overhead and are prone to hallucination artifacts that compromise physical fidelity.
Multi-exposure High Dynamic Range (HDR) imaging fuses Low Dynamic Range (LDR) images of varying exposures to recover scene radiance, but suffers from motion-induced misalignment in dynamic scenes. Recent Deep Neural Networks (DNNs), including Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) [yan2019attention, wu2018deep, yan2020deep, liu2022ghost, kalantari2017deep, yan2023unified, tel2023alignment], have greatly improved HDR reconstruction in dynamic scenes. Despite their advancements, these methods simply learn a mapping network between LDR-HDR paired data using pixel-wise losses [wang2021deep], which often yields perceptually unsatisfactory results. This limitation is particularly pronounced when essential content is missing in severely overexposed regions due to object or camera motion.
Recent advances in Latent Diffusion models (LDMs) [ho2020denoising, song2020denoising, dhariwal2021diffusion], particularly large-scale pre-trained text-to-image (T2I) models  [rombach2022high, podell2023sdxl, saharia2022photorealistic] trained on billions of image-text pairs, possess powerful natural image priors. These priors make LDMs a promising solution for addressing the perceptual shortcomings of current HDR reconstruction methods, especially in mitigating motion-induced ghosting and other complex artifacts. However, directly applying LDMs to high-quality HDR image generation remains non-trivial due to three key challenges: (1) Limited Dynamic Range Representation: The autoencoders of pre-trained LDMs compress linear HDR distributions into an 8-bit perceptual latent space, limiting their capacity to represent the dynamic range of HDR content [wang2025lediff, guan2025hdr]. (2) High Inference Cost: LDMs demand a substantial number of iteration steps in the denoising process, which is time-consuming even with a high-end GPU card [yan2023toward, hu2024generating]. (3) Content Hallucination: The generative nature of LDMs can lead to hallucinated content that deviates from realistic scenes. As shown in Fig. 1, UltraFusion [chen2025ultrafusion] produces human-perceptually consistent HDR-like content with a compressed dynamic range but suffer from high latency and hallucination.
These limitations motivate a paradigm shift in leveraging LDMs for HDR reconstruction. Our work draws from three key insights: First, inspired by the observation that an HDR image can be decomposed into a LDR component and a 8 bit-depth Gain Map (GM) [apple2021edr, adobe2024gainmap, google2024ultrahdr] that encodes the extended dynamic range, we reframe HDR reconstruction as a conditionally guided GM estimation task to fully exploit pre-trained LDM priors without domain mismatches. Second, since LDR inputs already provide rich structural and semantic cues, we initialize the denoising process from an informative prior (e.g., a regression-based GM estimate) rather than pure noise, significantly accelerating inference. Third, recognizing that regression-based models excel in content fidelity while LDMs favor perceptual quality, we leverage regression-based priors to guide the LDM’s denoising and latent decoding processes.
Building on these principles, we propose GMODiff, a gain map-driven one-step diffusion model for multi-exposure HDR reconstruction. Rather than training a LDM to directly reconstruct full HDR content, we adopt a two-stage fine-tuning strategy to adapt a pre-trained multi-step LDM for single-step GM refinement. Specifically, in the first stage, we train a Degradation-aware Regressor Network (DaReg) via a dual-learning strategy to produce an initial GM and extract spatial embeddings that highlight regions where the estimate is unreliable. This joint learning scheme provides the LDM with an informative prior for initialization and directs its attention to ambiguous or error-prone areas, significantly accelerating convergence and improving structural fidelity. In the second stage, we fine-tune the LDM to perform one-step denoising conditioned on these regression-based priors, enabling accurate GM correction and high-fidelity HDR reconstruction at dramatically reduced computational cost. In summary, our main contributions are threefold:
  • We propose GMODiff, a one-step diffusion framework that reframes HDR reconstruction as a conditionally guided gain map refinement task, fully leveraging pre-trained LDMs without redesigning the latent space.
  • We design a unified prior-guided refinement pipeline that seamlessly integrates regression-based fidelity with diffusion-driven perceptual enhancement: the DAReg produces an initial gain map and degradation-aware embeddings, which guide both the denoising process and the latent decoding of the LDM, effectively suppressing hallucinations while preserving structural accuracy.
  • Extensive experiments demonstrate that the proposed GMODiff delivers superior visual quality with significantly reduced ghosting and artifacts, while achieving high inference efficiency.
Figure 2: Overview of the proposed GMODiff framework. (a) We train a DaReg via a dual-learning strategy to produce two regression-based priors: an initial gain map $\hat{G}$ and spatial embeddings $c_{L}$. The $c_{L}$ implicitly encode regions where the $\hat{G}$ is unreliable, serving as a degradation-aware guidance prior. (b) One-Step Diffusion Refinement initializes the denoising process from $\hat{G}$ and performs single-step denoising conditioned on $c_{L}$ to generate a high-fidelity GM latent code $Z_{H}$. (c) The DA Decoder leverages encoder features from the initial gain map $\hat{G}$, guided by $c_{L}$, to recover fine image details while avoiding the introduction of artifacts inherent in $\hat{G}$.
2 Related Work
2.1 Multi-exposure HDR Reconstruction
Multi-exposure HDR imaging has advanced significantly, with methods broadly divided into traditional and deep learning-based categories. Traditional methods address motion artifacts via motion rejection [Gallo2009Artifact], registration [bogoni2000extending, kang2003high], and patch matching [sen2012robust, Hu2013deghosting], but their effectiveness hinges on preprocessing accuracy, often failing with large/complex motion. In contrast, state-of-the-art HDR reconstruction methods now predominantly leverage the powerful nonlinear modeling capabilities of DNNs. Recent advances incorporate attention mechanisms [yan2023unified, chen2022attention, yan2019attention] and ViT architectures [tel2023alignment, yan2023unified, liu2022ghost] to improve feature correspondence and capture long-range dependencies. Meanwhile, optical flow–guided approaches [kalantari2017deep, kong2024safnet] explicitly address misalignment by warping input frames according to estimated motion fields. Moreover, generative models, including Generative Adversarial Networks (GANs) [niu2021hdr] and diffusion models [yan2023toward, hu2024generating, chen2025ultrafusion, wang2025lediff], have been successfully adapted to synthesize photorealistic HDR content, even from imperfect or severely degraded inputs. Additional innovative frameworks, such as [li2025afunet, ni2025semantic, guan2025hdr, liaolearning], have further expanded the scope of DNN applications in this field, introducing novel architectures and training strategies to improve reconstruction quality. Despite these advances, most existing methods either learn a direct LDR-to-HDR mapping using pixel-wise losses, or train generative models from scratch on limited datasets. These often yield unsatisfactory results, especially when critical content is missing in severely over/underexposed regions due to motion.
2.2 Diffusion Models
Diffusion models have emerged as a powerful class of generative frameworks, capable of synthesizing high-fidelity data through iterative denoising of random noise. They have already demonstrated impressive performance across a wide range of low-level vision tasks—such as super-resolution [yue2025arbitrary, chen2025adversarial], denoising [du2025superpc], image enhancement [lan2025efficient], colorization [du2025superpc], and deblurring [xie2025diffusion], thanks to their ability to model fine-grained details and generate visually compelling outputs, often surpassing traditional image restoration methods. Despite their success, the practical deployment of diffusion models remains hindered by their high computational cost, which arises from both complex network architectures [song2020denoising, lu2022dpm]] and the requirement for hundreds of iterative sampling steps during inference. To mitigate this bottleneck, recent efforts have focused on accelerating diffusion by drastically reducing the number of denoising steps. However, aggressive step reduction typically leads to a noticeable degradation in output quality. Recently, img2img [parmar2024one] proposed a general framework that adapts pre-trained single-step diffusion models [xu2024ufogen, wang2023prolificdreamer] to new tasks and domains via adversarial learning. This approach enables efficient inference while leveraging the rich internal knowledge embedded in large-scale pre-trained diffusion models. Nevertheless, directly applying this efficient diffusion paradigm to HDR generation remains challenging. Most pre-trained diffusion models rely on autoencoders trained on 8-bit LDR images, which compress HDR’s high bit-depth into a perceptually encoded, low-dimensional latent space, severely limiting their ability to capture the full dynamic range of the real world.
Figure 3: (a) Limitations of the vanilla LDMs in generating HDR content. The limitation of the LDM VAE in encoding and decoding an HDR image, visualized in multiple exposure levels, which reveals a significant fidelity loss, especially in the shadow. (b) The GM encodes pixel-wise dynamic range adjustments for the LDR image and can be used to reconstruct the HDR image via element-wise multiplication.
3 Methodology
As shown in Fig. 2, GMODiff is trained in two stages. During the first stage (Sec. 3.2), we pretrain a Degradation-aware Regressor Network (DaReg) to predict an initial gain map and degradation-aware embeddings from the LDR inputs. During the second stage (Sec. 3.3), we fine-tune a pre-trained LDM using these regression priors to refine the initial gain map and reconstruct a high-quality HDR image.
3.1 Preliminaries
As shown in Fig. 3 (a), while the variational autoencoder (VAE) of a LDM can encode visually plausible LDR representations that resemble HDR content under standard viewing conditions, it fundamentally fails to preserve true high dynamic range—particularly in highlights and shadows. Moreover, since the denoising network operates in an 8-bit perceptual latent space, a naive solution to enable HDR generation would be to fine-tune both the VAE and the denoising network on HDR data. However, this approach demands large-scale HDR training sets and discards the rich generative priors learned from massive LDR images [wang2025lediff].
Recent advances in display technology have popularized dual-layer HDR formats [apple2021edr, google2024ultrahdr], which decompose an HDR image $\mathbf{H}$ into an LDR base layer $\mathbf{L}$ and an 8-bit gain map $\mathbf{G}$. The gain map is obtained by pixel-wise division of $\mathbf{H}$ by $\mathbf{L}$, followed by logarithmic compression. As shown in Fig. 3 (b), the original HDR image can then be reconstructed as:
$\mathbf{H}=(\mathbf{L}+\alpha)\cdot\exp_{2}\big(1+\mathbf{G}\cdot Q_{\max}\big),$ (1)
where $\mathbf{G}\in[0,1]$ is the normalized gain map, $\exp_{2}(\cdot)$ denotes the inverse log transform, $Q_{max}$ corresponds to the maximum gain value, and $\alpha=1/64$ is a scaling constant that preserves detail in near-black regions. Crucially, both LDR image and gain map are 8-bit depth, allowing us to directly exploit pre-trained LDMs without VAE retraining or latent space redesign.
3.2 Degradation-aware Regressor Network
The first stage produces two regression-based priors: an initial gain map $\hat{\mathbf{G}}$ and multi-scale spatial embeddings $\mathbf{c}_{L}$. These priors are used to initialize and condition the one-step diffusion refinement, ensuring high structural fidelity while enabling perceptual enhancement.
Overall Architecture. Our DaReg adopts a NAFNet backbone [chen2022simple] augmented with a lightweight degradation-aware segmentation head. Given the input LDR sequence $\{L_{1},L_{2},\ldots,L_{N}\}$, we first apply gamma correction to each $L_{i}$ to obtain its HDR version $H_{i}$, and form a six-channel tensor $X_{i}=[L_{i},H_{i}]$ by channel-wise concatenation. Following [yan2019attention, liu2022ghost, 10288540], an implicit alignment module [yan2019attention] processes the LDR inputs and fuses multi-exposure features before feeding them into the network. As shown in Fig. 2, DaReg encodes these features into multi-scale spatial embeddings:
$\mathbf{c}_{L}=\left\{\mathbf{c}_{L_{k}}\in\mathbb{R}^{d}\right\}_{k=1}^{K},$ (2)
where $\mathbf{c}_{L}$ aggregates encoder features across scales and flattens them into vector representations. The decoder reconstructs the initial gain map using $\mathbf{c}_{L}$, while the segmentation head predicts a binary reliability mask for degraded regions (e.g., due to motion or saturation). This mask endows $\mathbf{c}_{L}$ with awareness of uncertain areas, guiding subsequent refinement.
Initial GM Reconstruction. Given an input LDR sequence, the encoder extracts multi-scale features, where each stage consists of multiple Nonlinear Activation-Free Blocks [chen2022simple]. The channel dimensions at the four scales are set to 64, 128, 256, and 512, respectively. A symmetric decoder then reconstructs the initial gain map $\hat{\mathbf{G}}$ as: $\hat{\mathbf{G}}=D\left(\mathbf{c}_{L};\phi\right)$, where each decoder stage receives skip-connected features from the corresponding encoder stage to progressively reconstruct the initial GM.
[…]
$\mathcal{L}_{\text{gm}}=\|\hat{\mathbf{G}}-\mathbf{G}\|_{1}.$ (3)
Additionally, we enforce consistency between the reconstructed HDR image $\hat{\mathbf{H}}$ (from $\hat{\mathbf{G}}$) and the ground-truth $\mathbf{H}$:
$\mathcal{L}_{r}=\left\|\mathcal{T}(\mathbf{H})-\mathcal{T}(\hat{\mathbf{H}})\right\|_{1}+\lambda\left\|\phi_{i,j}(\mathcal{T}(\mathbf{H}))-\phi_{i,j}(\mathcal{T}(\hat{\mathbf{H}}))\right\|_{1},$ (4)
where $\mathcal{T}(\cdot)$ denotes the $\mu$-law funtion, $\phi_{i,j}(\cdot)$ is the $j$-th feature map of VGG19 after the $i$-th max-pooling layer, and $\lambda=1\mathrm{e}{-2}$ is a weighting hyperparameter.
Degradation-Aware Embedding. By incorporating a reconstruction objective, DaReg implicitly learns regression-based priors that facilitate dynamic range expansion. To better guide the LDM in refining the initial GM, we explicitly train it to identify a degradation-aware mask, such that $\mathbf{c}_{L}$ implicitly encodes regions where $\hat{\mathbf{G}}$ is unreliable. This guides the refinement process to focus on uncertain regions while avoiding hallucinating implausible details in well-reconstructed areas.
Specifically, given an input LDR sequence, we extract multi-scale latent representations with the DaReg encoder as before. For the degradation-aware mask $M$, we upsample all feature maps to the original resolution, concatenate them, and feed the aggregated features into a $1\times 1$ convolutional layer to produce $\bar{\mathbf{c}}_{L}$. The segmentation head $h(\cdot;\phi)$ processes $\bar{\mathbf{c}}_{L}$ through two convolutional layers with 64 hidden channels followed by a sigmoid activation, outputs a predict mask: $\hat{M}=h(\bar{\mathbf{c}}_{L};\phi),$ where $\hat{M}\in[0,1]$ indicates the predicted degradation at pixel $p$. We obtain the ground-truth mask $M$ using a simple yet effective pixel-consistency criterion:
$M(p)=\begin{cases}1&\text{if }\big\|\mathcal{T}(\mathbf{H}(p))-\mathcal{T}(\hat{\mathbf{H}}(p))\big\|_{1}>\sigma,\\ 0&\text{otherwise},\end{cases}$ (5)
where $p$ denotes a pixel location, $\|\cdot\|_{1}$ averages the absolute difference across RGB channels to yield a scalar, and $\sigma=4/255$ is a fixed threshold. This formulation is motivated by the observation that static regions yield consistent reconstructions—even with simple image blending—whereas motion or overexposure causes significant pixel-level discrepancies. The loss function adopted is the binary cross-entropy (BCE) loss between $\hat{M}$ and M{M}.
3.3 Prior-Guided One-Step Diffusion Refinement
In the second stage, our goal is to efficiently leverage the strong image priors of pre-trained LDMs to address degradations in the initial GM (e.g., ghosting and artifacts), thereby reconstructing high-quality HDR images.
[…]
$\hat{\mathbf{Z}}_{H}=F(\mathbf{Z}_{L};\mathbf{c})\triangleq\frac{\mathbf{Z}_{L}-\sqrt{1-\bar{\alpha}_{T_{L}}}\varepsilon_{\theta}\left(\mathbf{Z}_{L};\mathbf{c},T\right)}{\sqrt{\bar{\alpha}_{T}}},$ (6)
where $\varepsilon_{\theta}$ denotes the denoising network, $c$ is the guidance condition, $T\in[0,T]$ is the predefined diffusion timestep, and $\bar{\alpha}$ represents the predefined noise schedule. Specifically, we append trainable LoRA [hu2022lora] layers to the encoder $E_{\phi}$ and the diffusion network $\epsilon_{\phi}$, fine-tuning them into $E_{\theta}$ and $\epsilon_{\theta}$ using our training data. The condition $\mathbf{c}$ includes the regression prior $\mathbf{c}_{L}$ and the underexposed LDR image $L_{0}$. $\mathbf{c}_{L}$ is injected into the hidden layers of the denoising network, whereas $L_{0}$ is first encoded into a feature map, then concatenated with $\mathbf{Z}_{L}$ and fed into the denoising network, with the number of channels in the first convolutional layer adjusted accordingly.
[…]
Specifically, we obtain multi-scale features $\mathbf{Z}_{L}^{i}$ of the initial GM $\hat{\mathbf{G}}$ via the VAE encoder through forward propagation, while $\mathbf{\hat{Z}}_{H}$ is derived from a one-step diffusion process. Although the initial gain map, reconstructed via a regression approach, endows $\mathbf{Z}_{L}^{i}$ with rich detail fidelity, it may introduce noticeable artifacts. To address this, we propose a Degradation-Aware Refinement Block guided by the embedding $\mathbf{c}_{L}$, which leverages degradation-aware embedings to identify unreliable regions in $\mathbf{Z}_{L}^{i}$. As illustrated in Fig. 2 (c), the Degradation-Aware Refinement Block first concatenates the decoder intermediate features $\mathbf{\hat{Z}}_{H}^{i}$ with $\mathbf{Z}_{L}^{i}$ and applies a $3\times 3$ convolution to obtain refined features $\hat{\mathbf{Z}}^{i}$:
$\hat{\mathbf{Z}}^{i}=\operatorname{Conv}_{3\times 3}\big([\mathbf{\hat{Z}}_{H}^{i},\mathbf{Z}_{L}^{i}]\big),$ (7)
where $\cdot$] denotes channel-wise concatenation. The $\mathbf{c}_{L}$ is then passed through an MLP followed by a sigmoid activation to produce a channel-wise modulation vector:
$\mathbf{f}_{i}=\operatorname{Sigmoid}\big(\operatorname{MLP}(\mathbf{c}_{L})\big),$ (8)
$\mathbf{\hat{Z}}_{H}^{i+1}=\operatorname{Conv}_{3\times 3}\big(\hat{\mathbf{Z}}^{i}\odot\mathbf{f}_{i}\big)+\mathbf{\hat{Z}}_{H}^{i},$
where $\odot$ denotes channel-wise multiplication, and $\mathbf{\hat{Z}}_{H}^{i+1}$ represents the updated decoder feature after detail modulation and residual connection.
Training Strategy During the training phase, the loss function comprises the GM reconstruction loss and HDR reconstruction loss. Following [xu2024ufogen], we also compute the adversarial loss $\mathcal{L}_{\text{adv}}$ between the reconstructed HDR image and the GT HDR image in the $\mu$-law domain. The total objective for decoder training is:
$\mathcal{L}_{diff}=\mathcal{L}_{gm}+\mathcal{L}_{r}+\lambda\mathcal{L}_{adv}$ (9)
where $\lambda$ denotes the balancing parameter, which is empirically set to 0.001.
Additionally, considering that the initial GM $\hat{\mathbf{G}}$ obtained in the first stage may overfit to the training set, which impairs GMODiff’s ability to handle its degradations, we propose a group training strategy for GMODiff to learn robust degradation mitigation. Specifically, we divide the training set into N groups and train $N+1$ DaRegs accordingly, with the extra model trained on the entire training dataset. During GMODiff training, we randomly select one DaReg to provide $\hat{\mathbf{G}}$ and $\mathbf{c}_{L}$. The training objective for GMODiff is ultimately defined as the expected loss over such random selections:
$\mathcal{L}=\mathbb{E}{k\sim\text{Uniform}(0,N)}\left[\mathcal{L}_{diff}\left(\hat{G}_{k},\mathbf{c}_{L,k};H;G\right)\right]$ (10)
where $k$ is the index of a randomly selection.
4 Experiments
Table 1: Quantitative comparison across methods. Higher ($\uparrow$) is better for PU-PSNR,PU-SSIM,MANIQA, MUSIQ, CLIPIQA; lower ($\downarrow$) is better for LPIPS, DISTS, and NIQE. Best and second-best results per metric are highlighted in red and blue, respectively.
Metric DeepHDR [wu2018deep] NHDRR [yan2020deep] AHDRNet [yan2019attention] HDRGAN [niu2021hdr] HDR-Trans [liu2022ghost] SCTNet [tel2023alignment] SAFNet [kong2024safnet] DiffHDR [yan2023toward] AFUNet [li2025afunet] Ours
[…]
Figure 4: Visual comparisons are conducted on testing data, focusing on zoomed-in local areas of the HDR images estimated by our method and the compared techniques. Our model demonstrates the ability to generate HDR images of superior quality.
Figure 5: Visual comparisons are conducted on testing data , focusing on zoomed-in local areas of the HDR images estimated by our method and the compared techniques. Our model demonstrates the ability to generate HDR images of superior quality.
4.1 Experimental Settings
Datasets. We collect publicly available real-world datasets commonly used for multi-exposure HDR reconstruction from multiple sources [kalantari2017deep, tel2023alignment, kong2024safnet], resulting in a total of 836 image pairs. Among these, 758 pairs are used for training and the remaining 78 for testing. Each sample consists of three LDR images captured with exposure levels of either {-2, 0, +2} or {-3, 0, +3} stops. Notably, Challenge123 [kong2024safnet] also provides static LDR captures for each scene, enabling the generation of five additional training samples per scene compared to the other datasets.
Implementation Details. We implement our model in PyTorch and leverage the accelerate library for mixed-precision training with FP16. In stage one, we optimize DaReg using the Adam optimizer with a learning rate of 2e-4. The training data is preprocessed by extracting $256\times 256$ patches with a stride of 128, and we use a batch size of 16. DaReg adopts the NAFNet architecture [chen2022simple], with the encoder and decoder comprising [2, 4, 4, 8] and [2, 2, 2, 2] NAFBlocks per stage. The model is trained for 200K iterations on 2 NVIDIA RTX 4090 GPUs. In stage 2, we fine-tune Stable Diffusion 2.1 [rombach2022high] using LoRa with rank 16. DaReg is frozen and serves as a fixed module. We employ the AdamW optimizer with a learning rate of 5e-5, weight decay of 1e-10, and a batch size of 12. This stage is trained for 300K iterations on 4 NVIDIA RTX 4090 GPUs.
Evaluation Metrics.To ensure fair and comprehensive comparison, we employ both perceptual and distortion-based image quality metrics. For full-reference perceptual evaluation, we use LPIPS and DISTS, which assess perceptual similarity via features from pre-trained deep networks. For no-reference assessment, we adopt MANIQA, MUSIQ, and CLIPIQA, all trained to predict human-perceived quality. In particular, as these metrics are not HDR-specific, we computed them in the tone-mapped domain using a standard $\mu$-law operator. For distortion-based metrics, we include PU-PSNR and PU-SSIM, specifically tailored for HDR images, through the official implementation [azimi2021pu21].
4.2 Comparison with SOTA Methods
Compared Methods. We compare GMODiff against state-of-the-art CNN- and Transformer-based methods specifically designed for multi-exposure HDR image reconstruction, including DeepHDR [wu2018deep], NHDRR [yan2020deep], AHDRNet [yan2019attention], SCTNet [tel2023alignment], and HDR-Trans [liu2022ghost]. In addition, we also evaluate against recent advanced approaches: the GAN-based HDRGAN [niu2021hdr], the diffusion-based DiffHDR [yan2023toward], the optical-flow-based SAFNet [kong2024safnet], and the deep-unfolding-based AFUNet [li2025afunet]. For fair comparison, we retrain all these methods on our collected dataset using their official open-source implementations.
Qualitative Results. We present several challenging cases. As shown in Fig. 4, severe camera motion prevents accurate pixel-wise alignment, leaving only semantic cues available to guide reconstruction. The left case of Fig. 5 involves simultaneous human motion and over-saturation, resulting in significant information loss in the input LDR sequence. In the right example of Fig. 5, although the reference frame contains rich visual content, camera motion introduces misalignment that negatively impacts reconstruction quality. It can be observed that existing methods, which directly learn an end-to-end mapping from LDR to HDR, consistently produce noticeable artifacts or ghosting in the aforementioned challenging regions. Although DiffHDR attempts to tackle the ill-posed nature of HDR reconstruction by leveraging the generative capacity of diffusion models, its limited training data prevents it from synthesizing visually plausible results. In contrast, our proposed GMODiff integrates available input cues with strong image priors from pre-trained LDMs to effectively inpaint artifact-prone regions, while avoiding unrealistic hallucinations, resulting in more accurate and perceptually faithful reconstructions.
Quantitative Results. Quantitative comparisons on publicly available datasets are summarized in Tab. 1. Our method consistently outperforms competing approaches across a diverse set of perceptual metrics. Specifically, it achieves notable gains in both full-reference metrics (e.g., LPIPS and DISTS) and no-reference metrics (e.g., MUSIQ, MANIQA, and CLIPIQA), demonstrating its superior ability to reconstruct visually pleasing and perceptually faithful HDR images. Compared to AFUNet [li2025afunet], the state-of-the-art deep unfold method, our approach attains competitive performance in distortion-based metrics (PU-PSNR and PU-SSIM) while achieving significantly better perceptual quality—highlighting that leveraging regression priors to guide pre-trained latent diffusion models enables an effective trade-off between fidelity and perceptual realism. Moreover, in contrast to diffusion-based alternatives such as DiffHDR [yan2023toward], our method fully exploits the rich image priors embedded in large-scale pre-trained LDMs. DiffHDR, trained only on limited in-domain data, lacks robust generic priors and struggles in complex real-world scenarios.
4.3 Ablation Studies
Effects of HDR Reconstruction Loss. During the training of GMODiff, we replace the logarithmic compression in Eq. (1), which is non-invertible at zero, with $\mu$-law companding to enable direct supervision using ground-truth HDR images. This modification allows us to reconstruct the HDR image directly from the estimated gain map and apply pixel-wise supervision in the linear HDR domain. As shown in Tab. 2, ablating this reconstruction loss leads to clear performance degradation: minor inaccuracies in gain map estimation are amplified during HDR reconstruction, resulting in reduced fidelity and validating its importance for stable and accurate HDR reconstruction.
Effects of Degradation-Aware Priors. The degradation-aware priors, derived from intermediate features of the first-stage regression branch that predicts a degradation mask, implicitly encode spatially varying artifact patterns. Although the mask itself is not used as direct conditioning, its underlying feature representation provides useful cues about input-specific degradation. As shown in Tab. 2, removing these priors leads to a consistent drop across all metrics-both perceptual (e.g., LPIPS, MUSIQ) and distortion-based (e.g., PU-PSNR). This demonstrates that even implicit access to degradation-related features benefits HDR reconstruction, enabling the diffusion model to better adapt to input artifacts and produce more accurate and perceptually plausible results.
Effects of Group Training Strategy. To prevent DaReg from overfitting to the training set, which could hinder GMODiff’s ability to effectively learn artifact removal, we propose a Group Training strategy. As illustrated in Fig. 6, without this strategy, GMODiff can still mitigate degradation in the initial gain map estimation and perform basic artifact correction. However, when Group Training is applied, its artifact removal capability is significantly enhanced, yielding results with nearly imperceptible artifacts. Tab. 2 shows improved perceptual scores with Group Training, while distortion metrics slightly decrease, indicating that Group Training achieves a favorable trade-off between fidelity and perceptual quality.
Effects of DA Decoder. As shown in Tab. 2, removing the DA Decoder leads to a significant drop across all evaluation metrics, especially reference-based ones. This degradation stems from the highly compressed latent space of LDMs, which struggles to preserve fine details. To mitigate this limitation, we inject high-fidelity detail features extracted from intermediate layers of the gain map encoder and selectively refine them using the regression prior obtained from the first stage. As illustrated in Fig. 6, the DA Decoder effectively enhances texture fidelity while suppressing artifacts and ghosting present in the initial gain map. This improvement underscores the critical role of the DA Decoder in recovering high-quality details during HDR reconstruction, thereby substantially boosting the perceptual quality of the final output.
Table 2: Ablation study of the each component. The best and second best results are colored with red and blue, respectively.
[…]
DaReg 41.67 0.0432 0.5803 59.77
w/o $\mathcal{L}_{r}$ 41.54 0.0415 0.5878 60.74
w/o Group Training 42.19 0.0353 0.5848 60.65
w/o $c_{L}$ 41.76 0.0389 0.5859 60.59
w/o DA Decoder 35.78 0.1076 0.5763 59.89
GMODiff (ours) 41.95 0.0346 0.5884 60.81
Table 3: Inference time comparison on a $1080\times 1920$ image.
Time (s) $\downarrow$ AHDR DiffHDR AFUNet SCTNet UltraFusion Ours
Inference Time 0.24s 13.52s 6.16s 5.66s 93.4s 0.89s
4.4 Inference Time Analysis
Tab. 3 compares inference latency on $1920\times 1080$ HDR reconstruction. Our method is overall second-fastest, slightly slower than the lightweight CNN-based AHDRNet [yan2019attention], and significantly faster than transformer-heavy models (AFUNet [li2025afunet], SCTNet [tel2023alignment]), which, despite using efficient window-based attention, incur high latency due to deep stacking of Transformer blocks and processing at high spatial resolutions. Among diffusion-based approaches, UltraFusion [chen2025ultrafusion] requires over a minute per image, owing to patch-wise processing with 50 denoising steps, optical flow alignment, and ControlNet overhead. DiffHDR [yan2023toward], despite using only 5 steps, operates in pixel space without latent compression, resulting in high memory bandwidth and slow inference. In contrast, our method leverages a pre-trained LDM that performs computation in a highly compressed latent space with a single denoising step, achieving the fastest speed among diffusion-based methods. Moreover, it is compatible with efficient inference frameworks such as xFormers, which further reduce latency through memory, efficient attention and kernel fusion, making it practical for real-world applications.
Figure 6: Qualitative results of our ablation study.
5 Conclusion
We presented GMODiff, a gain map-driven one-step latent diffusion framework for efficient and high-quality HDR reconstruction from multi-exposure inputs. By formulating HDR reconstruction as gain map refinement in a dual-layer HDR representation, our method can directly exploit pre-trained latent diffusion priors without retraining the VAE or designing HDR-specific latent spaces. A degradation-aware regressor provides a strong initialization and spatial degradation cues, while the one-step latent denoising and degradation-aware decoding jointly suppress ghosting and artifacts. Comprehensive experiments demonstrate that GMODiff outperforms state-of-the-art methods across both perceptual and distortion-based metrics, while significantly reducing inference latency.

### v2

@@
-GMODiff: One-Step Gain Map Refinement with Diffusion Priors for HDR Reconstruction
+GMODiff: One-Step Gain Map Refinement with Diffusion Priors for Efficient HDR Reconstruction
 Abstract
@@
 Building on these principles, we propose GMODiff, a gain map-driven one-step diffusion model for multi-exposure HDR reconstruction.
+To the best of our knowledge, GMODiff is the first LDM-based method designed for multi-exposure HDR reconstruction.
 Rather than training a LDM to directly reconstruct full HDR content, we adopt a two-stage fine-tuning strategy to adapt a pre-trained multi-step LDM for single-step GM refinement.
 Specifically, in the first stage, we train a Degradation-aware Regressor Network (DaReg) via a dual-learning strategy to produce an initial GM and extract spatial embeddings that highlight regions where the estimate is unreliable.
-This joint learning scheme provides the LDM with an informative prior for initialization and directs its attention to ambiguous or error-prone areas, significantly accelerating convergence and improving structural fidelity.
+This learning scheme provides the LDM with an informative prior for initialization and directs its attention to ambiguous or error-prone areas, significantly accelerating convergence and improving structural fidelity.
 In the second stage, we fine-tune the LDM to perform one-step denoising conditioned on these regression-based priors, enabling accurate GM correction and high-fidelity HDR reconstruction at dramatically reduced computational cost.
@@
   • Extensive experiments demonstrate that the proposed GMODiff delivers superior visual quality with significantly reduced ghosting and artifacts, while achieving high inference efficiency.
-Figure 2: Overview of the proposed GMODiff framework.
-(a) We train a DaReg via a dual-learning strategy to produce two regression-based priors: an initial gain map $\hat{G}$ and spatial embeddings $c_{L}$.
-The $c_{L}$ implicitly encode regions where the $\hat{G}$ is unreliable, serving as a degradation-aware guidance prior.
-(b) One-Step Diffusion Refinement initializes the denoising process from $\hat{G}$ and performs single-step denoising conditioned on $c_{L}$ to generate a high-fidelity GM latent code $Z_{H}$.
-(c) The DA Decoder leverages encoder features from the initial gain map $\hat{G}$, guided by $c_{L}$, to recover fine image details while avoiding the introduction of artifacts inherent in $\hat{G}$.
 2 Related Work
@@
 Diffusion models have emerged as a powerful class of generative frameworks, capable of synthesizing high-fidelity data through iterative denoising of random noise.
-They have already demonstrated impressive performance across a wide range of low-level vision tasks—such as super-resolution [yue2025arbitrary, chen2025adversarial], denoising [du2025superpc], image enhancement [lan2025efficient], colorization [du2025superpc], and deblurring [xie2025diffusion], thanks to their ability to model fine-grained details and generate visually compelling outputs, often surpassing traditional image restoration methods.
-Despite their success, the practical deployment of diffusion models remains hindered by their high computational cost, which arises from both complex network architectures [song2020denoising, lu2022dpm]] and the requirement for hundreds of iterative sampling steps during inference.
+They have already demonstrated impressive performance across a wide range of low-level vision tasks-such as super-resolution [yue2025arbitrary, chen2025adversarial], denoising [du2025superpc], image enhancement [lan2025efficient], colorization [du2025superpc], and deblurring [xie2025diffusion], thanks to their ability to model fine-grained details and generate visually compelling outputs, often surpassing traditional image restoration methods.
+Despite their success, the practical deployment of diffusion models remains hindered by their high computational cost, which arises from both complex network architectures [song2020denoising, lu2022dpm] and the requirement for hundreds of iterative sampling steps during inference.
 To mitigate this bottleneck, recent efforts have focused on accelerating diffusion by drastically reducing the number of denoising steps.
@@
 Most pre-trained diffusion models rely on autoencoders trained on 8-bit LDR images, which compress HDR’s high bit-depth into a perceptually encoded, low-dimensional latent space, severely limiting their ability to capture the full dynamic range of the real world.
-Figure 3: (a) Limitations of the vanilla LDMs in generating HDR content.
-The limitation of the LDM VAE in encoding and decoding an HDR image, visualized in multiple exposure levels, which reveals a significant fidelity loss, especially in the shadow.
-(b) The GM encodes pixel-wise dynamic range adjustments for the LDR image and can be used to reconstruct the HDR image via element-wise multiplication.
 3 Methodology
 As shown in Fig.
-2, GMODiff is trained in two stages.
+3, GMODiff is trained in two stages.
 During the first stage (Sec.
@@
 As shown in Fig.
-3 (a), while the variational autoencoder (VAE) of a LDM can encode visually plausible LDR representations that resemble HDR content under standard viewing conditions, it fundamentally fails to preserve true high dynamic range—particularly in highlights and shadows.
+2 (a), while the variational autoencoder (VAE) of a LDM can encode visually plausible LDR representations that resemble HDR content under standard viewing conditions, it fundamentally fails to preserve true high dynamic range-particularly in highlights and shadows.
 Moreover, since the denoising network operates in an 8-bit perceptual latent space, a naive solution to enable HDR generation would be to fine-tune both the VAE and the denoising network on HDR data.
@@
 Crucially, both LDR image and gain map are 8-bit depth, allowing us to directly exploit pre-trained LDMs without VAE retraining or latent space redesign.
+Figure 2: (a) Limitations of the vanilla LDMs in generating HDR content.
+The limitation of the LDM VAE in encoding and decoding an HDR image, visualized in multiple exposure levels, which reveals a significant fidelity loss, especially in the shadow.
+(b) The GM encodes pixel-wise dynamic range adjustments for the LDR image and can be used to reconstruct the HDR image via element-wise multiplication.
+Figure 3: Overview of the proposed GMODiff framework.
+(a) We train a DaReg via a dual-learning strategy to produce two regression-based priors: an initial gain map $\hat{G}$ and spatial embeddings $c_{L}$.
+The $c_{L}$ implicitly encode regions where the $\hat{G}$ is unreliable, serving as a degradation-aware guidance prior.
+(b) One-Step Diffusion Refinement initializes the denoising process from $\hat{G}$ and performs single-step denoising conditioned on $c_{L}$ to generate a high-fidelity GM latent code $Z_{H}$.
+(c) The DA Decoder leverages encoder features from the initial gain map $\hat{G}$, guided by $c_{L}$, to recover fine image details while avoiding the introduction of artifacts inherent in $\hat{G}$.
 3.2 Degradation-aware Regressor Network
@@
 As shown in Fig.
-2, DaReg encodes these features into multi-scale spatial embeddings:
+3, DaReg encodes these features into multi-scale spatial embeddings:
 $\mathbf{c}_{L}=\left\{\mathbf{c}_{L_{k}}\in\mathbb{R}^{d}\right\}_{k=1}^{K},$ (2)
@@
 The decoder reconstructs the initial gain map using $\mathbf{c}_{L}$, while the segmentation head predicts a binary reliability mask for degraded regions (e.g., due to motion or saturation).
-This mask endows $\mathbf{c}_{L}$ with awareness of uncertain areas, guiding subsequent refinement.
+This mask will be fused with $\mathbf{c}_{L}$ after the first stage of training, which endows $\mathbf{c}_{L}$ with awareness of uncertain regions and guides the subsequent refinement.
 Initial GM Reconstruction.
@@
 A symmetric decoder then reconstructs the initial gain map $\hat{\mathbf{G}}$ as: $\hat{\mathbf{G}}=D\left(\mathbf{c}_{L};\phi\right)$, where each decoder stage receives skip-connected features from the corresponding encoder stage to progressively reconstruct the initial GM.
-[…]
+During training, we compute the ground-truth gain map $\mathbf{G}$ using Eq.
+(1), where $\mathbf{L}$ is the reference frame of the LDR sequence.
+To enable direct supervision with HDR images and ensure differentiability at zero, we replace the logarithmic compression with the $\mu$-law function.
+Accordingly, the $\exp_{2}(\cdot)$ in Eq.
+(1) is substituted by the inverse $\mu$-law mapping: $R(x)=\frac{e^{x\cdot\log(1+\mu)}-1}{\mu}$, with $\mu$=100.
+We supervise the initial gain map prediction $\hat{\mathbf{G}}$ via an $\mathcal{L}_{1}$ regression loss:
 $\mathcal{L}_{\text{gm}}=\|\hat{\mathbf{G}}-\mathbf{G}\|_{1}.$ (3)
@@
 $\mathcal{L}_{r}=\left\|\mathcal{T}(\mathbf{H})-\mathcal{T}(\hat{\mathbf{H}})\right\|_{1}+\lambda\left\|\phi_{i,j}(\mathcal{T}(\mathbf{H}))-\phi_{i,j}(\mathcal{T}(\hat{\mathbf{H}}))\right\|_{1},$ (4)
-where $\mathcal{T}(\cdot)$ denotes the $\mu$-law funtion, $\phi_{i,j}(\cdot)$ is the $j$-th feature map of VGG19 after the $i$-th max-pooling layer, and $\lambda=1\mathrm{e}{-2}$ is a weighting hyperparameter.
+where $\mathcal{T}(\cdot)$ denotes the $\mu$-law function, $\phi_{i,j}(\cdot)$ is the $j$-th feature map of VGG19 after the $i$-th max-pooling layer, and $\lambda=1\mathrm{e}{-2}$ is a weighting hyperparameter.
 Degradation-Aware Embedding.
 By incorporating a reconstruction objective, DaReg implicitly learns regression-based priors that facilitate dynamic range expansion.
-To better guide the LDM in refining the initial GM, we explicitly train it to identify a degradation-aware mask, such that $\mathbf{c}_{L}$ implicitly encodes regions where $\hat{\mathbf{G}}$ is unreliable.
-This guides the refinement process to focus on uncertain regions while avoiding hallucinating implausible details in well-reconstructed areas.
+To better guide the LDM in refining the initial GM, we additionally introduce a segmentation head to predict a degradation-aware mask, which allows the refinement process to focus on uncertain regions and avoids hallucinating implausible details in well-reconstructed areas.
 Specifically, given an input LDR sequence, we extract multi-scale latent representations with the DaReg encoder as before.
-For the degradation-aware mask $M$, we upsample all feature maps to the original resolution, concatenate them, and feed the aggregated features into a $1\times 1$ convolutional layer to produce $\bar{\mathbf{c}}_{L}$.
-The segmentation head $h(\cdot;\phi)$ processes $\bar{\mathbf{c}}_{L}$ through two convolutional layers with 64 hidden channels followed by a sigmoid activation, outputs a predict mask: $\hat{M}=h(\bar{\mathbf{c}}_{L};\phi),$ where $\hat{M}\in[0,1]$ indicates the predicted degradation at pixel $p$.
+For the degradation-aware mask $M$, we process the encoder features to the same number of channels using $1\times 1$ convolutions, and upsample them to the original resolution for concatenation to produce $\bar{\mathbf{c}}_{L}$.
+The segmentation head $h(\cdot;\phi)$ processes $\bar{\mathbf{c}}_{L}$ through three convolutional layers with 64 hidden channels, followed by a sigmoid activation, and outputs a predict mask: $\hat{M}=h(\bar{\mathbf{c}}_{L};\phi),$ where $\hat{M}\in[0,1]$ denotes the predicted degradation probability at each pixel $p$.
+Notably, to prevent the gradient of the segmentation head from affecting the reconstruction accuracy, we perform a detach operation before feature processing.
+The obtained degradation-aware mask is then compressed through a lightweight CNN network in the next training stage and fused with the original $\mathbf{c}_{L}$ to serve as the conditional guidance for the LDMs.
 We obtain the ground-truth mask $M$ using a simple yet effective pixel-consistency criterion:
@@
 where $p$ denotes a pixel location, $\|\cdot\|_{1}$ averages the absolute difference across RGB channels to yield a scalar, and $\sigma=4/255$ is a fixed threshold.
-This formulation is motivated by the observation that static regions yield consistent reconstructions—even with simple image blending—whereas motion or overexposure causes significant pixel-level discrepancies.
+This formulation is motivated by the observation that static regions yield consistent reconstructions-even with simple image blending-whereas motion or overexposure causes significant pixel-level discrepancies.
 The loss function adopted is the binary cross-entropy (BCE) loss between $\hat{M}$ and M{M}.
@@
 Although the initial gain map, reconstructed via a regression approach, endows $\mathbf{Z}_{L}^{i}$ with rich detail fidelity, it may introduce noticeable artifacts.
-To address this, we propose a Degradation-Aware Refinement Block guided by the embedding $\mathbf{c}_{L}$, which leverages degradation-aware embedings to identify unreliable regions in $\mathbf{Z}_{L}^{i}$.
+To address this, we propose a Degradation-Aware Refinement Block guided by the embedding $\mathbf{c}_{L}$, which leverages degradation-aware embeddings to identify unreliable regions in $\mathbf{Z}_{L}^{i}$.
 As illustrated in Fig.
@@
 Training Strategy During the training phase, the loss function comprises the GM reconstruction loss and HDR reconstruction loss.
-Following [xu2024ufogen], we also compute the adversarial loss $\mathcal{L}_{\text{adv}}$ between the reconstructed HDR image and the GT HDR image in the $\mu$-law domain.
+Following [DBLP:journals/corr/abs-2503-06537, guo2025compression, xu2024ufogen], we also compute the adversarial loss $\mathcal{L}_{\text{adv}}$ between the predicted latent gain map $\mathbf{\hat{Z}}_{H}$ and the GT latent gain map $\mathbf{{Z}}_{H}$, where the discriminator consists of a frozen pre-trained denoising UNet encoder and two trainable MLP layers.
 The total objective for decoder training is:
@@
 During GMODiff training, we randomly select one DaReg to provide $\hat{\mathbf{G}}$ and $\mathbf{c}_{L}$.
-The training objective for GMODiff is ultimately defined as the expected loss over such random selections:
+The training objective for GMODiff is ultimately defined as the expected loss over such a random selections:
 $\mathcal{L}=\mathbb{E}{k\sim\text{Uniform}(0,N)}\left[\mathcal{L}_{diff}\left(\hat{G}_{k},\mathbf{c}_{L,k};H;G\right)\right]$ (10)
@@
 4 Experiments
-Table 1: Quantitative comparison across methods.
-Higher ($\uparrow$) is better for PU-PSNR,PU-SSIM,MANIQA, MUSIQ, CLIPIQA; lower ($\downarrow$) is better for LPIPS, DISTS, and NIQE.
-Best and second-best results per metric are highlighted in red and blue, respectively.
-Metric DeepHDR [wu2018deep] NHDRR [yan2020deep] AHDRNet [yan2019attention] HDRGAN [niu2021hdr] HDR-Trans [liu2022ghost] SCTNet [tel2023alignment] SAFNet [kong2024safnet] DiffHDR [yan2023toward] AFUNet [li2025afunet] Ours
+Table 1: Quantitative comparison on the in-distribution test split of our collected dataset. Best and second-best results per metric are highlighted in red and blue, respectively.
+Metric DeepHDR [wu2018deep] AHDRNet [yan2019attention] SAFNet [niu2021hdr] HDR-Trans [liu2022ghost] SCTNet [tel2023alignment] AFUNet [li2025afunet] LE-Diff [wang2025lediff] GMDD [guan2025hdr] DiffHDR [yan2023toward] Ours
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 2: Quantitative comparisons on the Prabhakar et al.’s [prabhakar2019fast] dataset. Best and second-best results per metric are highlighted in red and blue, respectively.
+Metric DeepHDR [wu2018deep] AHDRNet [yan2019attention] SAFNet [niu2021hdr] HDR-Trans [liu2022ghost] SCTNet [tel2023alignment] AFUNet [li2025afunet] LE-Diff [wang2025lediff] GMDD [guan2025hdr] DiffHDR [yan2023toward] Ours
 […]
@@
 Datasets.
-We collect publicly available real-world datasets commonly used for multi-exposure HDR reconstruction from multiple sources [kalantari2017deep, tel2023alignment, kong2024safnet], resulting in a total of 836 image pairs.
-Among these, 758 pairs are used for training and the remaining 78 for testing.
-Each sample consists of three LDR images captured with exposure levels of either {-2, 0, +2} or {-3, 0, +3} stops.
-Notably, Challenge123 [kong2024safnet] also provides static LDR captures for each scene, enabling the generation of five additional training samples per scene compared to the other datasets.
+Scenes and motion variations within a single dataset are limited.
+To better validate the model’s generalization ability under diverse real-world conditions, we collect publicly available real-world datasets [kong2024safnet, tel2023alignment, kalantari2017deep] commonly used for multi-exposure HDR reconstruction, which cover a wide range of illumination and motion conditions.
+With data augmentation and cropping, we obtain over 400k training pairs in total.
+For evaluation, we conduct quantitative comparisons on datasets with ground truth, including the test split of our training data (in-distribution) and the Prabhakar [prabhakar2019fast] dataset (out-of-distribution, 562 scenes) for real-world robustness testing.
+Furthermore, we adopt the datasets of Sen et al.
+[sen2012robust] and Tursun et al.
+[Tursun2016data] for qualitative visual assessment only, as they lack ground truth.
 Implementation Details.
@@
 In stage one, we optimize DaReg using the Adam optimizer with a learning rate of 2e-4.
-The training data is preprocessed by extracting $256\times 256$ patches with a stride of 128, and we use a batch size of 16.
+The training data is preprocessed by extracting $256\times 256$ patches, and we use a batch size of 16.
 DaReg adopts the NAFNet architecture [chen2022simple], with the encoder and decoder comprising [2, 4, 4, 8] and [2, 2, 2, 2] NAFBlocks per stage.
-The model is trained for 200K iterations on 2 NVIDIA RTX 4090 GPUs.
+The model is trained for 200K iterations on 8 NVIDIA RTX 4090 GPUs.
+After training, DaReg is frozen, and the segmentation head is independently trained for 5K iterations.
 In stage 2, we fine-tune Stable Diffusion 2.1 [rombach2022high] using LoRa with rank 16.
-DaReg is frozen and serves as a fixed module.
+DaReg and the segmentation head are frozen and serve as fixed modules.
 We employ the AdamW optimizer with a learning rate of 5e-5, weight decay of 1e-10, and a batch size of 12.
 This stage is trained for 300K iterations on 4 NVIDIA RTX 4090 GPUs.
-Evaluation Metrics.To ensure fair and comprehensive comparison, we employ both perceptual and distortion-based image quality metrics.
-For full-reference perceptual evaluation, we use LPIPS and DISTS, which assess perceptual similarity via features from pre-trained deep networks.
+Evaluation Metrics.
+To ensure fair and comprehensive comparison, we employ both perceptual and distortion-based image quality metrics.
+For full-reference perceptual evaluation, we use DISTS, which assess perceptual similarity via features from pre-trained deep networks.
 For no-reference assessment, we adopt MANIQA, MUSIQ, and CLIPIQA, all trained to predict human-perceived quality.
 In particular, as these metrics are not HDR-specific, we computed them in the tone-mapped domain using a standard $\mu$-law operator.
-For distortion-based metrics, we include PU-PSNR and PU-SSIM, specifically tailored for HDR images, through the official implementation [azimi2021pu21].
+For distortion-based metrics, we include PU-PSNR and PU-SSIM, specifically tailored for HDR images, through the official implementation [azimi2021pu21] with a peak luminance of 1000 nits.
+We compute HDR-VDP2 on linear HDR images with the “rgb-native” color encoding.
+The data is treated as display-referred, scaled to a peak luminance of 1000 nits, with a viewing angle of $30^{\circ}$ and viewing distance of 0.55 m.
 4.2 Comparison with SOTA Methods
 Compared Methods.
-We compare GMODiff against state-of-the-art CNN- and Transformer-based methods specifically designed for multi-exposure HDR image reconstruction, including DeepHDR [wu2018deep], NHDRR [yan2020deep], AHDRNet [yan2019attention], SCTNet [tel2023alignment], and HDR-Trans [liu2022ghost].
-In addition, we also evaluate against recent advanced approaches: the GAN-based HDRGAN [niu2021hdr], the diffusion-based DiffHDR [yan2023toward], the optical-flow-based SAFNet [kong2024safnet], and the deep-unfolding-based AFUNet [li2025afunet].
+We compare GMODiff against state-of-the-art CNN- and Transformer-based methods specifically designed for multi-exposure HDR image reconstruction, including DeepHDR [wu2018deep], AHDRNet [yan2019attention], SCTNet [tel2023alignment], and HDR-Trans [liu2022ghost].
+In addition, we also evaluate against recent advanced approaches: the diffusion-based DiffHDR [yan2023toward], the optical-flow-based SAFNet [kong2024safnet], and the deep-unfolding-based AFUNet [li2025afunet].
 For fair comparison, we retrain all these methods on our collected dataset using their official open-source implementations.
+We further compare with existing LDM-based HDR reconstruction methods [google2024ultrahdr, guan2025hdr, wang2025lediff], which leverage pre-trained LDMs and are fine-tuned on significantly larger datasets than ours.
+Notably, since UltraFusion [google2024ultrahdr] is a multi-exposure fusion method that cannot generate HDR images, we only include it for qualitative comparison.
+Figure 6: Qualitative comparison of HDR reconstruction results with state-of-the-art LDM-based methods.
+Our GMODiff outperforms existing LDM-based methods in recovering fine details, preserving physical illumination consistency and suppressing artifacts in overexposed/motion regions.
 Qualitative Results.
-We present several challenging cases.
+We present several challenging scenarios.
 As shown in Fig.
-4, severe camera motion prevents accurate pixel-wise alignment, leaving only semantic cues available to guide reconstruction.
-The left case of Fig.
-5 involves simultaneous human motion and over-saturation, resulting in significant information loss in the input LDR sequence.
-In the right example of Fig.
-5, although the reference frame contains rich visual content, camera motion introduces misalignment that negatively impacts reconstruction quality.
-It can be observed that existing methods, which directly learn an end-to-end mapping from LDR to HDR, consistently produce noticeable artifacts or ghosting in the aforementioned challenging regions.
-Although DiffHDR attempts to tackle the ill-posed nature of HDR reconstruction by leveraging the generative capacity of diffusion models, its limited training data prevents it from synthesizing visually plausible results.
-In contrast, our proposed GMODiff integrates available input cues with strong image priors from pre-trained LDMs to effectively inpaint artifact-prone regions, while avoiding unrealistic hallucinations, resulting in more accurate and perceptually faithful reconstructions.
+4, severe camera motion invalidates accurate pixel-wise alignment, leaving only semantic cues for reconstruction.
+The left case in Fig.
+5 involves both human motion and over-saturation, causing severe information loss in the input LDR sequence.
+In the right example, despite the informative reference frame, camera motion still induces misalignment and degrades reconstruction quality.
+As shown in Fig.
+4 and 5, existing end-to-end methods consistently suffer from obvious artifacts or ghosting in these challenging regions.
+Although DiffHDR exploits the generative ability of diffusion models, it still fails to produce visually plausible results due to limited training data.
+In contrast, our GMODiff combines input cues with powerful priors from pre-trained LDMs to robustly inpaint artifact-prone regions without unrealistic hallucinations, yielding more accurate and perceptually faithful results.
+We further compare GMODiff with existing LDM-based methods in Fig.
+6.
+GMDD [guan2025hdr] and LE-Diff [wang2025lediff] exhibit severe detail loss because LDMs operate in a compressed latent space.
+Specifically, GMDD cannot recover missing content in over-exposed areas, while LE-Diff fails to model physical illumination, leading to large brightness and color deviations from the GT.
+Ultrafusion [google2024ultrahdr] is trained solely on LDR images and thus cannot generate real HDR outputs; it also relies heavily on optical-flow alignment, producing obvious artifacts when alignment fails.
+By contrast, guided by the regression module, GMODiff strictly follows physical constraints and avoids latent-compression information loss via the DA Decoder, generating more realistic and visually pleasing results.
+Qualitative comparisons with Prabhakar et al. [prabhakar2019fast], Sen et al. [sen2012robust], and Tursun et al. [Tursun2016data] are provided in the supplementary material.
 Quantitative Results.
 Quantitative comparisons on publicly available datasets are summarized in Tab.
-1.
+1 and Tab.
+2.
 Our method consistently outperforms competing approaches across a diverse set of perceptual metrics.
-Specifically, it achieves notable gains in both full-reference metrics (e.g., LPIPS and DISTS) and no-reference metrics (e.g., MUSIQ, MANIQA, and CLIPIQA), demonstrating its superior ability to reconstruct visually pleasing and perceptually faithful HDR images.
-Compared to AFUNet [li2025afunet], the state-of-the-art deep unfold method, our approach attains competitive performance in distortion-based metrics (PU-PSNR and PU-SSIM) while achieving significantly better perceptual quality—highlighting that leveraging regression priors to guide pre-trained latent diffusion models enables an effective trade-off between fidelity and perceptual realism.
+Specifically, it achieves notable gains in both full-reference metrics and no-reference metrics, demonstrating its superior ability to reconstruct visually pleasing and perceptually faithful HDR images.
+Compared to AFUNet [li2025afunet], the state-of-the-art deep unfold method, our approach attains competitive performance in distortion-based metrics (PU-PSNR and PU-SSIM) while achieving significantly better perceptual quality-highlighting that leveraging regression priors to guide pre-trained latent diffusion models enables an effective trade-off between fidelity and perceptual realism.
 Moreover, in contrast to diffusion-based alternatives such as DiffHDR [yan2023toward], our method fully exploits the rich image priors embedded in large-scale pre-trained LDMs.
 DiffHDR, trained only on limited in-domain data, lacks robust generic priors and struggles in complex real-world scenarios.
+Notably, since LE-Diff [wang2025lediff] recalculates physical brightness during the reconstruction process, its HDR results suffer from severe brightness discrepancies, leading to extremely low full-reference metrics.
+Therefore, we do not report it in the table.
+The user study is shown in Figure 9, with details provided in the supplementary material.
 4.3 Ablation Studies
-Effects of HDR Reconstruction Loss.
-During the training of GMODiff, we replace the logarithmic compression in Eq.
-(1), which is non-invertible at zero, with $\mu$-law companding to enable direct supervision using ground-truth HDR images.
-This modification allows us to reconstruct the HDR image directly from the estimated gain map and apply pixel-wise supervision in the linear HDR domain.
-As shown in Tab.
-2, ablating this reconstruction loss leads to clear performance degradation: minor inaccuracies in gain map estimation are amplified during HDR reconstruction, resulting in reduced fidelity and validating its importance for stable and accurate HDR reconstruction.
-Effects of Degradation-Aware Priors.
-The degradation-aware priors, derived from intermediate features of the first-stage regression branch that predicts a degradation mask, implicitly encode spatially varying artifact patterns.
-Although the mask itself is not used as direct conditioning, its underlying feature representation provides useful cues about input-specific degradation.
-As shown in Tab.
-2, removing these priors leads to a consistent drop across all metrics-both perceptual (e.g., LPIPS, MUSIQ) and distortion-based (e.g., PU-PSNR).
-This demonstrates that even implicit access to degradation-related features benefits HDR reconstruction, enabling the diffusion model to better adapt to input artifacts and produce more accurate and perceptually plausible results.
-Effects of Group Training Strategy.
-To prevent DaReg from overfitting to the training set, which could hinder GMODiff’s ability to effectively learn artifact removal, we propose a Group Training strategy.
-As illustrated in Fig.
-6, without this strategy, GMODiff can still mitigate degradation in the initial gain map estimation and perform basic artifact correction.
-However, when Group Training is applied, its artifact removal capability is significantly enhanced, yielding results with nearly imperceptible artifacts.
-Tab.
-2 shows improved perceptual scores with Group Training, while distortion metrics slightly decrease, indicating that Group Training achieves a favorable trade-off between fidelity and perceptual quality.
-Effects of DA Decoder.
-As shown in Tab.
-2, removing the DA Decoder leads to a significant drop across all evaluation metrics, especially reference-based ones.
-This degradation stems from the highly compressed latent space of LDMs, which struggles to preserve fine details.
-To mitigate this limitation, we inject high-fidelity detail features extracted from intermediate layers of the gain map encoder and selectively refine them using the regression prior obtained from the first stage.
-As illustrated in Fig.
-6, the DA Decoder effectively enhances texture fidelity while suppressing artifacts and ghosting present in the initial gain map.
-This improvement underscores the critical role of the DA Decoder in recovering high-quality details during HDR reconstruction, thereby substantially boosting the perceptual quality of the final output.
-Table 2: Ablation study of the each component. The best and second best results are colored with red and blue, respectively.
+To validate the effectiveness of each core component in the proposed GMODiff framework, we conduct comprehensive ablation experiments on the benchmark test set, with quantitative results reported in Table LABEL:abb and qualitative results presented in Fig.
+7.
+We take the full GMODiff model as the baseline to ablate key modules one by one, and also evaluate the standalone performance of DaReg for comparison.
+DaReg-Seg refers to the joint training of the reconstruction branch and segmentation head, which leads to a noticeable performance drop.
+In the second stage of the framework, removing the DA Decoder causes a significant performance decline, as the latent space of LDM fails to preserve the rich detail information contained in the initial GM.
+Thanks to the rational design of the overall framework, the model only experiences slight performance degradation when other modules are ablated.
+Compared with the original regression-based DaReg model, GMODiff achieves a favorable trade-off between perceptual quality and distortion metrics with acceptable computational overhead.
+Figure 7: Qualitative results of our ablation study.
+Table 3: Inference time comparison on a $1080\times 1920$ image.
+Time (s) $\downarrow$ AHDR [yan2019attention] DiffHDR [yan2023toward] AFUNet [kong2024safnet] SCTNet [tel2023alignment] UltraFusion [chen2025ultrafusion] GMDD [guan2025hdr] LE-Diff [wang2025lediff] Ours
+Inference Time 0.24s 13.52s 6.16s 5.66s 93.4s 39.2s 132.7s 0.93s
+Figure 8: Ablation study of the each component. The best and second best results are colored with red and blue.
 […]
-DaReg 41.67 0.0432 0.5803 59.77
-w/o $\mathcal{L}_{r}$ 41.54 0.0415 0.5878 60.74
-w/o Group Training 42.19 0.0353 0.5848 60.65
-w/o $c_{L}$ 41.76 0.0389 0.5859 60.59
-w/o DA Decoder 35.78 0.1076 0.5763 59.89
-GMODiff (ours) 41.95 0.0346 0.5884 60.81
-Table 3: Inference time comparison on a $1080\times 1920$ image.
-Time (s) $\downarrow$ AHDR DiffHDR AFUNet SCTNet UltraFusion Ours
-Inference Time 0.24s 13.52s 6.16s 5.66s 93.4s 0.89s
+DaReg 36.87 0.5819 60.57
+DaReg-Seg 35.47 0.5803 59.77
+w/o $\mathcal{L}_{r}$ 36.60 0.5875 60.84
+w/o Group Training 36.76 0.5858 60.75
+w/o DA Mask 36.49 0.5862 60.73
+w/o DA Decoder 30.57 0.5754 59.93
+GMODiff (ours) 36.68 0.5881 60.96
+Figure 9: The user study results on the benchmark test set. Details can be found in the supplementary materials.
 4.4 Inference Time Analysis
@@
 DiffHDR [yan2023toward], despite using only 5 steps, operates in pixel space without latent compression, resulting in high memory bandwidth and slow inference.
+Since LE-Diff [wang2025lediff] contains two LDMs that separately predict results for different exposures, it takes more than 2 minutes to infer one complete HDR image.
 In contrast, our method leverages a pre-trained LDM that performs computation in a highly compressed latent space with a single denoising step, achieving the fastest speed among diffusion-based methods.
 Moreover, it is compatible with efficient inference frameworks such as xFormers, which further reduce latency through memory, efficient attention and kernel fusion, making it practical for real-world applications.
-Figure 6: Qualitative results of our ablation study.
 5 Conclusion